(1/x) - [1/(x+4)]=1/3
i know i need to multiply by 3x(x+4) but i dont know what to do after that

Answer by fractalier(6550)

Yes, you're on the right track...from
(1/x) - [1/(x+4)] = 1/3, we do multiply by the LCD 3x(x+4)...we get
3x(x+4)[(1/x) - [1/(x+4)] = 1/3]
Now cancel out what cancels term by term...
3(x+4) - 3x = x(x+4)
3x + 12 - 3x = x^2 + 4x
x^2 + 4x - 12 = 0
now factor and solve
(x + 6)(x - 2) = 0
x = -6 and x = 2 (check to see that they're "legal" solutions, which they are)